A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. Combining 44 ounces of Solution A with 66 ounces of Solution B gives 110 ounces of a mixture that is 55% salt. Combining 14 ounces of Solution A with 56 ounces of Solution B gives 70 ounces of a mixture that is 60 salt. What are the percentages of salt in the two solutions?
Answer by stanbon(75887)

A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. Combining 44 ounces of Solution A with 66 ounces of Solution B gives 110 ounces of a mixture that is 55% salt. Combining 14 ounces of Solution A with 56 ounces of Solution B gives 70 ounces of a mixture that is 60 salt. What are the percentages of salt in the two solutions?
----
a is % of salt in A
b is % of salt in B
------
Note: salt + salt = salt
------
44a + 66b = 55*110
14a + 56b = 60*70
-------
Modify:
4a + 6b = 55*10
4a + 16b = 60*20
--------------------
Subtract and solve for "b":
10b = 650
b = 65%
---
Solve for "a":
4a + 16*65 = 60*20
4a = 160
a = 40% (
==============
Cheers,
Stan H.
========================